After settlement, the average weekly wage in a factory had increased from Rs. 8 to 12 and standard deviation had increased from Rs. 1 to 1.5. The wages have become higher and more uniform, after settlement. Comment.


1
Expert's answer
2021-08-24T08:12:41-0400

Coefficient of Variation "=" "\left(\frac{Standard\:Deviation}{Mean}\right)\times 100\%"


Earlier Average weekly wages "=8"

Mean "=8"

Standard Deviation "=1"

Coefficient of Variation"=\left(\:\frac{1}{8}\right)\times 100\:\:=\:12.5\:\%"


After settlement, Average weekly wages = 12

Mean = 12

Standard Deviation = 1.5

Coefficient of Variation"=\left(\frac{1.5}{12}\right)\times 100\:\:=\:12.5\:\%"


"12>8"

Hence wages have become higher is Correct

Coefficient of Variation is same

Hence no impact on uniformity wages have become more uniform is not correct


Wages have become higher is Correct but more uniform is not correct